# Experimental Analysis of Biasing Voltage Generation in Wave-Controlled RIS

### v1

@@ III-B Biasing Transmission Line Design
The BTL design is that of a printed meander (or serpentine) transmission line, chosen for its compact size, phase delay control, and ease of fabrication; see Fig. 5. It is made of copper and implemented on a grounded dielectric substrate RO3010 with relative permittivity $\epsilon_{r}=11.2$, $\tan\delta=0.0022$, and thickness $h=0.64$ mm. The BTL board consists of $27$ identical unit cells, where the two cells at the ends have been modified to include SMA connectors that allow connection of the signal generator and integration of possible external loads, if required, to ensure optimal performance within the desired frequency range. The proposed design details and geometry are displayed in Fig. 5. The dimensions of the BTL unit cell are (in mm): $d_{x}=20$, $C=5$, $E=7.64$, $F=11$, $G=1.8$, $H=50$, $L_{\mathrm{p}}=131.42$, $W=2.6$.
Each unit cell of the BTL has a path length $L_{\mathrm{p}}$ and a spatial period $d_{x}$ along the $x$-axis (the same as for the RIS board), and is connected to a rectifier that extracts the desired dc voltage to bias the varactor located on the RIS board. The BTL signal travels along the microstrip with phase velocity $c/n_{{\mathrm{eff}}}$ where $n_{{\mathrm{eff}}}=\sqrt{\epsilon}_{{\mathrm{eff}}}=2.94$ is the effective refractive index of the microstrip and $\epsilon_{{\mathrm{eff}}}=8.66$ is its effective permittivity, which is approximately calculated as [20]
@@
and is considered dispersionless in the frequency range of operation. The characteristic impedance of the microstrip is $Z_{0}=19.23$ $\Omega$. In this case, the slowness factor, used in (1)-(4), is also determined by the path length $L_{\mathrm{p}}$ and the period $d_{x}$, and is given by $n_{{\mathrm{slow}}}=n_{{\mathrm{geom}}}n_{{\mathrm{eff}}}$, where $n_{{\mathrm{geom}}}=L_{\mathrm{p}}/d_{x}=6.57$ represents the geometric index that allows for projection of the standing wave pattern from the printed meander transmission line onto the $x$-axis. Thus, the resulting slowness factor is $n_{{\mathrm{slow}}}=19.34$. Although the segments $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ of the BTL follow the same meander line geometry as the main section $L$, these two additional segments can have a different geometry as they are not connected to any rectifier. However, the values of $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ must be normalized with respect to the geometric index in order to be used in (1)-(4). Therefore, the total length of the BTL along the $x$-axis is $L_{\mathrm{tot}}=540$ mm, with $L=520$ mm, and $L_{\mathrm{left}}=L_{\mathrm{right}}=0.5d_{x}=10$ mm, while the total path length of the meander microstrip is $3548.3$ mm.
@@
Using printed meander microstrip topology for the BTL design offers the significant advantage of increasing the slowness of the waves in the $x$-direction, allowing the use of low frequencies for the BSWs. This advantageous feature helps mitigate unwanted interference from the BTL harmonics on the RF system. The scattering parameters have been analyzed to assess the losses and reflection performance of the BTL design. The selected frequency range spans from 0 to 300 MHz, which is much lower than the RF band (here, around 2.45 GHz). When the RIS is used with several standing waves, as in [13, 16], this design allows for a fundamental frequency of approximately 7.2 MHz. This ensures that at least 21 modes are available below 300 MHz, which is a sufficient number of degrees of freedom for RIS control. The frequency $f_{\mathrm{b}}$ of the fundamental BSW decreases when the number $M$ of RIS elements increases. Fig. 6 presents the transmission coefficient of the BTL design (without considering the losses of the 27 rectifiers), obtained from a full-wave simulation in CST Studio Suite. The results show a gradual decrease with frequency due to ohmic and dielectric losses, reaching a minimum value of 2.44-2.44 dB at 300 MHz. Although not depicted in the figure, the reflection coefficient with Port 2 matched to its characteristic impedance remains consistently below 34-34 dB across the entire frequency range, indicating minimal mismatch. The BSWs are formed along the BTL due to an open or short termination at $x=-L_{\mathrm{left}}$.
@@
Figure 5: Geometry of the BTL made of a meander microstrip line over a grounded dielectric layer (not shown). The rectifiers of the ac BSW are at the bottom of the figure, one per RIS element. The voltages $w(x_{m},t)$, m=0,1,,M1m=0,1,...,M-1 are obtained from the lower portion of the BTL. The node of the rectified voltage $w(x_{m})$ is subsequently connected to the coplanar waveguide on the RIS board (connection not shown here).
@@ IV-A Theoretical Analysis and Experimental Results
To attain the full capabilities of the wave-controlled RIS, the maximum voltage of $w(x_{m})$ measured after the rectifiers must be high enough to enable the widest possible reflection phase range. This voltage amplitude is influenced by the proper selection of the frequency $f_{\mathrm{b}}$ and the BTL termination, either short- or open-circuited. For a short-circuited termination at $x=-L_{\mathrm{left}}$, the input impedance at $x=L+L_{\mathrm{right}}$ is given by $Z_{\mathrm{in}}=jZ_{0}\mathrm{tan}\left(\kappa\right)$, where $Z_{0}$ is the characteristic impedance of the microstrip and $\kappa=(\omega_{\mathrm{b}}n_{\mathrm{slow}}/c)L_{\mathrm{tot}}$. In the particular case where $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$, with $n=1,2,3,$..., and $f_{\mathrm{b,0}}$ is chosen such that the wavelength $\lambda_{\mathrm{b,0}}=c/(n_{\mathrm{slow}}f_{\mathrm{b,0}})=4L_{\mathrm{tot}}$, the input impedance is $Z_{\mathrm{in},n}=jZ_{0}\mathrm{tan}\left(n\pi/2\right)$. Under this condition, $Z_{\mathrm{in},n}$ becomes infinite for odd multiples of $f_{\mathrm{b,0}}$ (i.e., $n=1,3,5,$...), and zero for even multiples (i.e., $n=2,4,6,$...). As a result, the voltage at $x=L+L_{\mathrm{right}}$ reaches a maximum for odd multiples and a minimum for even multiples of $f_{\mathrm{b,0}}$.
@@
Figure 8: Printed meander BTL backed by the grounded dielectric substrate RO3010 connected to the Siglent SDG6052X AWG, and the RIGOL DP832 dc power supply.
@@
In order to experimentally compare the voltage distribution along the BTL for even and odd multiples of $f_{\mathrm{b,0}}$, we use the experimental setup shown in Fig. 8, which includes an AWG, power supply, decoupling elements, and BTL. For the BTL design presented here, $f_{\mathrm{b,0}}=7.18$ MHz. After setting $V_{\textrm{g}}=10$ V, Fig. 9 shows the voltages $w(x_{m})$ measured with a multimeter at positions $x_{m}=md_{x}$, $m=1,2,3,$..., after each rectifier along the BTL terminated in a short circuit, for six BSWs with frequencies that are multiples of $f_{\mathrm{b,0}}$. The results reveal a maximum amplitude difference exceeding $3$ V between odd and even multiples of $f_{\mathrm{b,0}}$. However, the minimum values remain close to $W_{0}$ in all cases, slightly falling below $W_{0}$ due to losses associated with the rectifiers. The voltages at the first and last elements show slight deviations from the expected values, with these deviations increasing at higher frequencies since the measurements were taken half a period $(d_{x}/2)$ before the ends of the BTL, corresponding to the lengths of $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$.
@@
Figure 9: Measured dc bias voltage $w(x_{m})$ along the BTL at positions $x_{m}=md_{x}$ for six individual BSWs, with the BTL terminated by a short circuit at $x=-L_{\mathrm{left}}$. The red and blue curves represent cases where $f_{\mathrm{b}}$ is an even or odd multiple of $f_{\mathrm{b,0}}$, respectively.
@@
Similarly, for the case where the BTL is terminated in an open circuit at $x=-L_{\mathrm{left}}$, the input impedance at $x=L+L_{\mathrm{right}}$ is $Z_{\mathrm{in}}=jZ_{0}\mathrm{cot}\left(\kappa\right)$. By choosing $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$ as before, the input impedance for $n$-multiples of $f_{\mathrm{b,0}}$ becomes $Z_{\mathrm{in},n}=jZ_{0}\mathrm{cot}\left(n\pi/2\right)$. This causes $Z_{\mathrm{in},n}$ to become infinite for even multiples of $f_{\mathrm{b,0}}$ and zero for odd multiples, resulting in a maximum voltage at $x=L+L_{\mathrm{right}}$ for odd multiples and a minimum for even multiples of $f_{\mathrm{b,0}}$. Fig. 10 presents a comparison of the voltages $w(x_{m})$ measured along the BTL with an open-circuit termination. Similar to previous cases, the results reveal a maximum amplitude difference of over $3$ V between odd and even multiples of $f_{\mathrm{b,0}}$, with higher amplitudes for even multiples in this case. As before, the minimum values remain close to $W_{0}$ for all cases, while slight deviations are observed in the first and last elements due to the $d_{x}/2$ shift before the BTL endpoints.
@@
The difference observed in the amplitude for open- and short-circuited terminations is attributed to the effect of the output impedance $Z_{\mathrm{g}}$ of the AWG that injects the periodic time-domain signal at $nf_{\mathrm{b,0}}$. As an example, Fig. 11 shows the maximum dc biasing voltage $w_{\mathrm{max}}=\max\limits_{m}\ w(x_{m})=W_{\mathrm{b}}+W_{\mathrm{0}}$ versus the frequency $f_{\mathrm{b}}$ for a few values of the AWG impedance $Z_{\mathrm{g}}$ from $Z_{0}$ ($19.23$ $\Omega$, red curve) to $100$ $\Omega$ (green curve), assuming a short-circuit termination. These results are obtained using an analytical model of the BTL.
@@
In summary, from a practical perspective, two different design criteria can be employed to maximize $w_{\mathrm{max}}$ and ensure that it remains constant for all frequencies. The first approach involves selecting $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$ with $n$ odd, which guarantees maximization regardless of the relationship between $Z_{\mathrm{g}}$ and $Z_{0}$. For the second approach, the BTL is designed with $Z_{0}=Z_{\mathrm{g}}$, which ensures maximization of $w(x_{m})$ for every value of $f_{\mathrm{b}}$. The second approach offers an additional advantage when multiple BSWs are used [16], as it effectively doubles the number of biasing modes within the same frequency range while also making $f_{\mathrm{b}}$ independent of the total length $L_{\mathrm{tot}}$.
@@ IV-B Comparison with Numerical Model
Fig. 13 compares the voltage $w(x_{m})$ evaluated using the ADS model with that from experimental results for two different BSW frequencies: one using an even multiple of $f_{\mathrm{b,0}}$ and the other using an odd multiple, both with a short-circuit termination. The results show good agreement between simulation and measurement in the case of the even multiple of $f_{\mathrm{b,0}}$, where $w_{\mathrm{max}}$ is minimized. The discrepancy between simulation and measurement increases with distance from the short-circuit termination. This can be attributed to the fact that the TL segments that model the BTL in ADS are straight and therefore do not account for geometric effects that are neglected in the model, such as the curvature present in the meanderline design, as well as coupling effects between parallel segments.
@@
Fig. 14 compares the reflection-phase distribution along the RIS, determined using the relationship in Fig. 4(4(a)) using both the ADS and measured voltages. As expected, the case with the even multiple $2f_{\mathrm{b,0}}$ exhibits good agreement between simulation and measurement, although the phase variation is limited, ranging from approximately $150\degree$ to $45\degree$ due to the relatively narrow range of biasing voltages. In contrast, the odd multiple case, $5f_{\mathrm{b,0}}$, shows a much broader phase variation, ranging from about $150^{\circ}$ to $120^{\circ}$, but with a more noticeable mismatch between ADS and experimental results for large $m$. Finally, Fig. 15 illustrates the expected scattered radiation patterns calculated using the field reflection magnitude and phase for each RIS element obtained from both the simulated and measured cases. The scattered pattern is evaluated using the normalized array factor, defined as
@@
where $\alpha_{m}$ and $R_{m}$, previously shown in Fig. 4(4(a)) , represent the phase and magnitude of the reflection coefficient for the $m$-th unit cell given the applied voltage $w(x_{m})$ as determined by (4). For the two selected values $2f_{\mathrm{b,0}}$ and $5f_{\mathrm{b,0}}$, the resulting radiation patterns concentrate the reflected energy near $\theta=0^{\circ}$ and $\theta=\pm 32^{\circ}$. However, the key takeaway is the comparison between the two cases, where, as expected, the even multiple of $f_{\mathrm{b,0}}$ yields better agreement.

### v2

@@ III-B Biasing Transmission Line Design
The goal is to control the RIS reflection coefficients independently using the BSWs $w_{n}(x,t)$ in the BTL. After rectification at $x_{m}=md_{x}$, control of the dc biasing voltages $w(m)$ leads to control of the reflection magnitudes and phases $R_{m}$ and $\alpha_{m}$ on each cell.
The BTL is a printed microstrip meandered line (or serpentine), chosen for its compact size, phase delay control, and ease of fabrication; see Fig. 5. It is made of copper and implemented on a grounded dielectric substrate RO3010 with relative permittivity $\epsilon_{r}=11.2$, $\tan\delta=0.0022$, and thickness $h=0.64$ mm. The BTL board consists of $27$ identical unit cells, where the two cells at the ends have been modified to include SMA connectors that allow connection of the signal generator and integration of possible external loads, if required, to ensure optimal performance within the desired frequency range. The proposed design details and geometry are displayed in Fig. 5. The dimensions of the BTL unit cell are (in mm): $d_{x}=20$, $C=5$, $E=7.64$, $F=11$, $G=1.8$, $H=50$, $L_{\mathrm{p}}=131.42$, $W=2.6$.
Each unit cell of the BTL has a curvilinear path length $L_{\mathrm{p}}$ and a spatial period $d_{x}$ along the $x$ direction (the same as for the RIS board), and is connected to a rectifier that extracts the desired dc voltage to bias the varactor located on the RIS board. The BTL signal travels along the microstrip with phase velocity $c/n_{{\mathrm{eff}}}$ where $n_{{\mathrm{eff}}}=\sqrt{\epsilon}_{{\mathrm{eff}}}=2.94$ is the effective refractive index of the microstrip and $\epsilon_{{\mathrm{eff}}}=8.66$ is its effective permittivity, which is approximately calculated as [23]
@@
and is considered dispersionless in the frequency range of operation. The characteristic impedance of the microstrip is $Z_{0}=19.23$ $\Omega$. In this case, the slowness factor, used in (1)-(4), is also determined by the path length $L_{\mathrm{p}}$ and the period $d_{x}$, and is given by $n_{{\mathrm{slow}}}=n_{{\mathrm{geom}}}n_{{\mathrm{eff}}}$, where $n_{{\mathrm{geom}}}=L_{\mathrm{p}}/d_{x}=6.57$ represents the geometric index that allows for projection of the standing wave pattern from the printed meandered line onto the $x$ direction. Thus, the resulting slowness factor is $n_{{\mathrm{slow}}}=19.34$. Although the segments $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ of the BTL follow the same meandered line geometry as the main section $L$, these two additional segments can have a different geometry as they are not connected to any rectifier. However, the values of $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$ must be normalized with respect to the geometric index in order to be used in (1)-(4). Therefore, the total length of the BTL along the $x$ direction is $L_{\mathrm{tot}}=540$ mm, with $L=520$ mm, and $L_{\mathrm{left}}=L_{\mathrm{right}}=0.5d_{x}=10$ mm, while the total path length of the meandered microstrip is $3548.3$ mm.
@@
Using printed meandered microstrip topology for the BTL design offers the significant advantage of increasing the slowness of the waves in the $x$ direction, allowing the use of low frequencies for the BSWs. This advantageous feature helps mitigate unwanted interference from the BTL harmonics on the RF system. The scattering parameters have been analyzed to assess the losses and reflection performance of the BTL design. The selected frequency range spans from 0 to 300 MHz, which is much lower than the RF band (here, around 2.45 GHz). When the RIS is used with several standing waves, as in [13, 16], this design allows for a fundamental frequency of approximately 7.2 MHz. This ensures that at least 21 modes are available below 300 MHz, which is a sufficient number of degrees of freedom for RIS control. The frequency $f_{\mathrm{b}}$ of the fundamental BSW decreases when the number $M$ of RIS elements increases. Fig. 6 presents the transmission coefficient of the BTL design (without considering the losses of the 27 rectifiers), obtained from a full-wave simulation in CST Studio Suite. The results show a gradual decrease with frequency due to ohmic and dielectric losses, reaching a minimum value of 2.44-2.44 dB at 300 MHz. Although not depicted in the figure, the reflection coefficient with Port 2 matched to its characteristic impedance remains consistently below 34-34 dB across the entire frequency range, indicating minimal mismatch. The BSWs are formed along the BTL due to an open or short termination at $x=-L_{\mathrm{left}}$.
@@
Figure 5: Geometry of the BTL formed by a meandered microstrip line over a grounded dielectric layer (not shown). The rectifiers of the ac BSW are at the bottom of the figure, one per RIS element. The voltages $w(x_{m},t)$, m=0,1,,M1m=0,1,...,M-1 are obtained from the lower portion of the BTL. The node of the rectified voltage $w(m)$ is subsequently connected to the coplanar waveguide on the RIS board (connection not shown here).
@@ IV-A Theoretical Analysis and Experimental Results
To attain the full capabilities of the wave-controlled RIS, the maximum value of the voltage $w(m)$ measured after the rectifiers must be high enough to enable the widest possible reflection phase range. This voltage amplitude is influenced by proper selection of the frequency $f_{\mathrm{b}}$ and the BTL termination, either short- or open-circuited. For a short-circuited termination at $x=-L_{\mathrm{left}}$, the input impedance at $x=L+L_{\mathrm{right}}$ is given by $Z_{\mathrm{in}}=jZ_{0}\mathrm{tan}\left(\kappa\right)$, where $Z_{0}$ is the characteristic impedance of the microstrip and $\kappa=(\omega_{\mathrm{b}}n_{\mathrm{slow}}/c)L_{\mathrm{tot}}$. In the particular case where $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$, with $n=1,2,3,$..., and $f_{\mathrm{b,0}}$ is chosen such that the wavelength $\lambda_{\mathrm{b,0}}=c/(n_{\mathrm{slow}}f_{\mathrm{b,0}})=4L_{\mathrm{tot}}$, the input impedance is $Z_{\mathrm{in},n}=jZ_{0}\mathrm{tan}\left(n\pi/2\right)$. Under this condition, $Z_{\mathrm{in},n}$ becomes infinite for odd multiples of $f_{\mathrm{b,0}}$ (i.e., $n=1,3,5,$...), and zero for even multiples (i.e., $n=2,4,6,$...). As a result, the voltage at $x=L+L_{\mathrm{right}}$ reaches a maximum for odd multiples and a minimum for even multiples of $f_{\mathrm{b,0}}$.
@@
Figure 8: Printed meandered BTL backed by the grounded dielectric substrate RO3010 connected to the Siglent SDG6052X AWG, and the RIGOL DP832 dc power supply.
@@
In order to experimentally compare the voltage distribution along the BTL for even and odd multiples of $f_{\mathrm{b,0}}$, we use the experimental setup shown in Fig. 8, which includes an AWG, power supply, decoupling elements, and BTL. For the BTL design presented here, $f_{\mathrm{b,0}}=7.18$ MHz. After setting $V_{\textrm{g}}=10$ V, Fig. 9 shows the voltages $w(m)$ measured with a multimeter at positions $x_{m}=md_{x}$, m=1,2,,M1m=1,2,...,M-1, after each rectifier along the BTL terminated in a short circuit, for six BSWs with frequencies that are multiples of $f_{\mathrm{b,0}}$. The results reveal a maximum amplitude difference exceeding $3$ V between odd and even multiples of $f_{\mathrm{b,0}}$. However, the minimum values remain close to $W_{0}$ in all cases, slightly falling below $W_{0}$ due to losses associated with the rectifiers. The voltages at the first and last elements show slight deviations from the expected values, with these deviations increasing at higher frequencies since the measurements were taken half a period $(d_{x}/2)$ before the ends of the BTL, corresponding to the lengths of $L_{\mathrm{left}}$ and $L_{\mathrm{right}}$.
@@
Figure 9: Measured dc bias voltage $w(m)$ along the BTL at positions $x_{m}=md_{x}$ for six individual BSWs, with the BTL terminated by a short circuit at $x=-L_{\mathrm{left}}$. The red and blue curves represent cases where $f_{\mathrm{b}}$ is an even or odd multiple of $f_{\mathrm{b,0}}$, respectively.
@@
Similarly, for the case where the BTL is terminated in an open circuit at $x=-L_{\mathrm{left}}$, the input impedance at $x=L+L_{\mathrm{right}}$ is $Z_{\mathrm{in}}=jZ_{0}\mathrm{cot}\left(\kappa\right)$. By choosing $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$ as before, the input impedance for $n$-multiples of $f_{\mathrm{b,0}}$ becomes $Z_{\mathrm{in},n}=jZ_{0}\mathrm{cot}\left(n\pi/2\right)$. This causes $Z_{\mathrm{in},n}$ to become infinite for even multiples of $f_{\mathrm{b,0}}$ and zero for odd multiples, resulting in a maximum voltage at $x=L+L_{\mathrm{right}}$ for odd multiples and a minimum for even multiples of $f_{\mathrm{b,0}}$. Fig. 10 presents a comparison of the voltages $w(m)$ measured along the BTL with an open-circuit termination. Similar to previous cases, the results reveal a maximum amplitude difference of over $3$ V between odd and even multiples of $f_{\mathrm{b,0}}$, with higher amplitudes for even multiples in this case. As before, the minimum values remain close to $W_{0}$ for all cases, while slight deviations are observed in the first and last elements due to the $d_{x}/2$ shift before the BTL endpoints.
@@
The difference observed in the amplitude for open- and short-circuited terminations is attributed to the effect of the output impedance $Z_{\mathrm{g}}$ of the AWG that injects the periodic time-domain signal at $nf_{\mathrm{b,0}}$. As an example, Fig. 11 shows the maximum dc biasing voltage $w_{\mathrm{max}}=\max\limits_{m}\ w(m)=W_{\mathrm{b}}+W_{\mathrm{0}}$ versus the frequency $f_{\mathrm{b}}$ for a few values of the AWG impedance $Z_{\mathrm{g}}$ from $Z_{0}$ ($19.23$ $\Omega$, red curve) to $100$ $\Omega$ (green curve), assuming a short-circuit termination. These results are obtained using an analytical model of the BTL.
@@
In summary, from a practical perspective, two different design criteria can be employed to maximize $w_{\mathrm{max}}$ and ensure that it remains constant for all frequencies. The first approach involves selecting $f_{\mathrm{b}}=nf_{\mathrm{b,0}}$ with $n$ odd, which guarantees maximization regardless of the relationship between $Z_{\mathrm{g}}$ and $Z_{0}$. For the second approach, the BTL is designed with $Z_{0}=Z_{\mathrm{g}}$, which ensures maximization of $w(m)$ for every value of $f_{\mathrm{b}}$. The second approach offers an additional advantage when multiple BSWs are used [16], as it effectively doubles the number of biasing modes within the same frequency range while also making $f_{\mathrm{b}}$ independent of the total length $L_{\mathrm{tot}}$.
@@ IV-B Comparison with Numerical Model
Fig. 13 compares the voltage $w(m)$ evaluated using the ADS model with that from experimental results for two different BSW frequencies: one using an even multiple of $f_{\mathrm{b,0}}$ and the other using an odd multiple, both with a short-circuit termination. The results show good agreement between simulation and measurement in the case of the even multiple of $f_{\mathrm{b,0}}$, where $w_{\mathrm{max}}$ is minimized. The discrepancy between simulation and measurement increases with distance from the short-circuit termination. This can be attributed to the fact that the TL segments that model the BTL in ADS are straight and therefore do not account for geometric effects that are neglected in the model, such as the curvature present in the meandered line design, as well as coupling effects between parallel segments.
@@
Fig. 14 compares the reflection-phase distribution along the RIS, determined using the relationship in Fig. 4(4(a)) using both ADS and measured voltages. As expected, the case with the even multiple $2f_{\mathrm{b,0}}$ exhibits good agreement between simulation and measurement, although the phase variation is limited, ranging from approximately $150^{\circ}$ to $45^{\circ}$ due to the relatively narrow range of biasing voltages. In contrast, the odd multiple case, $5f_{\mathrm{b,0}}$, shows a much broader phase variation, ranging from about $150^{\circ}$ to $120^{\circ}$, but with a more noticeable mismatch between ADS and experimental results for large $m$. Finally, Fig. 15 illustrates the expected scattered radiation patterns calculated using the field reflection magnitude and phase for each RIS element obtained from both the simulated and measured cases. The scattered pattern is evaluated using the normalized array factor, defined as
@@
where $\alpha_{m}$ and $R_{m}$, previously shown in Fig. 4(4(a)) , represent the phase and magnitude of the reflection coefficient for the $m$-th unit cell given the applied voltage $w(m)$ as determined by (4). For the two selected values $2f_{\mathrm{b,0}}$ and $5f_{\mathrm{b,0}}$, the resulting radiation patterns concentrate the reflected energy near $\theta=0^{\circ}$ and $\theta=\pm 32^{\circ}$. However, the key takeaway is the comparison between the two cases, where, as expected, the even multiple of $f_{\mathrm{b,0}}$ yields better agreement.